OpenStudy (anonymous):

expand using the binomial theorem: (4-5i)^3 Remember that i=sqrt -1

OpenStudy (anonymous):

(4-5i)^3 = -236-115i

OpenStudy (anonymous):

that's not what expanding is, but thanks for trying.

OpenStudy (jamesj):

By the binomial theorem: (a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3 So here \[(4-5i)^3 = 4^3 + 3.4^2(-5i) +3.4.(-5i)^2 + (-5i)^3\] \[ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = 64 + -240i + 300i^2 + -125i^3 \] \[ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = 64 + -240i -300 + 125i \] \[ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = -236 - 115i \]